What is the specific heat of silver if the temperature of a 15.4 g sample of silver is increased by 11.2 °C when 40.5 J of heat is added?
1
Expert's answer
2020-06-29T07:45:19-0400

C = Q / (m×DT )        C= Specific heat,  Energy(Q)= 40.5 J, Mass(m)= 15.4g,

C = 40.5 /(11.2×15.4) J/(g×°C)                                 Change of temperature (DT) = 11.2°C

= 0.234 J/(g×°C) (Answer)

The specific heat of silver is 0.234 J/(g×°C).